OpenStudy (gabylovesyou):

Solve the following system by using graph paper or graphing technology. 3x – 2y = 7 9x – 6y = 3 What is the solution to the system? No Solutions Infinite Solutions (3, 2) (3, –2)

OpenStudy (gabylovesyou):

@ash2326

OpenStudy (ash2326):

@Gabylovesyou Multiply first equation by 3

OpenStudy (gabylovesyou):

3x - 2y = 21 ?

OpenStudy (ash2326):

\[3\times (3x-2y)=3\times 7\] You have to multiply by 3 on both sides. We can't do any operation on only one side be it add, sub, mul or division

OpenStudy (gabylovesyou):

oh lol well i get 3 x (3x - 2y) = 21x

OpenStudy (ash2326):

Simplify it

OpenStudy (ash2326):

Right side is not x, but multiplication sign (x)

OpenStudy (gabylovesyou):

i simplify that ^^

OpenStudy (ash2326):

Simplify it and compare it with the second equation. What do you see?

OpenStudy (gabylovesyou):

pair of intersecting lines ?

OpenStudy (ash2326):

Don't arrive at the answer, tell me what do you see in the two equations? Is there any similarity?

OpenStudy (gabylovesyou):

x = 0? mmm idk help?

OpenStudy (ash2326):

Write the first equation here, after multiplying by 3 :) I'll help you figure this

OpenStudy (gabylovesyou):

3x (3x - 2y) = 3 x 7

OpenStudy (ash2326):

You have to multiply 3 to the left side and right side. 3x7=21 isn't it?????

OpenStudy (gabylovesyou):

yes lol

OpenStudy (ash2326):

@Gabylovesyou You have to multiply 3 to the left side also do it, would you?

OpenStudy (gabylovesyou):

.-. ... im lost now

OpenStudy (ash2326):

\[3\times (3x-2y)=21\] \[3\times 3x- 3\times 2y=21\] \[9x-6y=21\] Do you get this?

OpenStudy (gabylovesyou):

oh lol yeses

OpenStudy (ash2326):

Now write this and the second equation here

OpenStudy (gabylovesyou):

9x - 6y = 21 the second equation as in the solution for the variable y

OpenStudy (ash2326):

The second equation which is in the question

OpenStudy (gabylovesyou):

9x - 6y = 3

OpenStudy (ash2326):

what about the left side of both equations? Is it same?

OpenStudy (gabylovesyou):

yes

OpenStudy (ash2326):

so if you compare them, it means that \[21=3\] Is this true?

OpenStudy (gabylovesyou):

yes

OpenStudy (ash2326):

is 21 = 3?